A sum amounting to $3.35 consists of quarters and dimes. If there are 23 coins in all how many are quarters?
Answer by Maths68(1474)

Total Sum of coins = $3.35
Total Coins = 23
Dimes = x
Quarters = 23-x
Value of a Dime = $0.1
Value of a Quarter =$0.25
(value of a Dime)(Dimes)+(value of a Quarter)(Quarters)=Total Sum of Coins
(0.1)(x)+(0.25)(23-x)=3.35
0.1x+5.75-0.25x=3.35
0.1x-0.25x=3.35-5.75
-0.15x=-2.4
-0.15x/-0.15=-2.4/-0.15
x=16


Dimes = x = 16
Quarters = 23-x = 23-16 = 7
There are 7 Quarters


Check
======
(0.1)(16)+(7)(0.25)=3.35
1.6+1.75=3.35
3.35=3.35